(D^2+2D+5)y=xe^x


1
Expert's answer
2021-01-11T18:35:20-0500

Given: "(D^2+2D+5)y=xe^x"

Here, "D=\frac{d}{dx}"

So, we can rewrite it as: "\frac{d^2y}{dx^2}+2\frac{dy}{dx}+5y=xe^x" ... (i)

First, we find "y_c" , complimentary function.

Consider "\frac{d^2y}{dx^2}+2\frac{dy}{dx}+5y=0"

Its auxiliary equation is: "m^2+2m+5=0"

On solving by quadratic formula,

"m=\frac{-2\pm\sqrt{2^2-4(1)(5)}}{2(1)}=\frac{-2\pm\sqrt{4-20}}{2}=-1\pm2i"

Thus, the solution of homogeneous equation is:

"y_c=e^{-1x}(A\cos2x+B\sin2x)=Ae^{-x}\cos2x+Be^{-x}\sin2x"

Now, we find particular solution of non-homogeneous equation.

Assume "y=(ax+b)e^x" be the solution ...(ii)

We find a and b by substitution and comparison.

On differentiating (ii) w.r.t "x" ,

"\frac{dy}{dx}=(ax+b)e^x+(a)e^x
\\ \Rightarrow\frac{dy}{dx}=(ax+a+b)e^x"

Again, differentiating it w.r.t "x" ,

"\frac{d^2y}{dx^2}=(ax+a+b)e^x+ae^x
\\ \Rightarrow \frac{d^2y}{dx^2}=(ax+2a+b)e^x"

Now substituting these values in (i), we get,

"(ax+2a+b)e^x+2(ax+a+b)e^x+5(ax+b)e^x=xe^x
\\ \Rightarrow (ax+2a+b)+2(ax+a+b)+5(ax+b)=x"

"\Rightarrow (8a)x+(4a+8b)=1x+0"

On comparing,

"8a=1\Rightarrow a=\frac 1 8
\\ 4a+8b=0 \Rightarrow 8b=-4a \Rightarrow b=-\frac{a}{2}"

So, "b=-\frac1 2 \times \frac 1 8=-\frac 1 {16}"

So, "y_p=(\frac1 8x-\frac 1 {16})e^x=\frac{xe^x}{8}-\frac{e^x}{16}"

Thus, "y(x)=y_c+y_p"

"=Ae^{-x}\cos2x+Be^{-x}\sin2x+\frac{xe^x}{8}-\frac{e^x}{16}"

Answer: "y(x)=Ae^{-x}\cos2x+Be^{-x}\sin2x+\frac{xe^x}{8}-\frac{e^x}{16}"